MICO 22 nd February 2010 Terry Hart (MOM) - Decay Solenoid and Target - MICE Machine Physics runs - User Run Plans.
Published byModified over 5 years ago
Presentation on theme: "MICO 22 nd February 2010 Terry Hart (MOM) - Decay Solenoid and Target - MICE Machine Physics runs - User Run Plans."— Presentation transcript:
MICO 22 nd February 2010 Terry Hart (MOM) - Decay Solenoid and Target - MICE Machine Physics runs - User Run Plans
Decay Solenoid and Target 22 Feb, 20102 Decay Solenoid (from Matt Hills) Retightened a few small leaks in Swagelok joints Found leak in pressure regulator on coldbox – Explains why coldbox vacuum not as good as that of DS or compressor – Regulator designed to seal with internal pressure instead of internal vacuum – Leak shouldn’t be a problem when system is filled and pressurized Refilled system and restarted gas cleaning Cool down could start in about a day – Liquid nitrogen to continue gas cleaning arrived today – Need < 10 ppm impurities to start cooldown (now at around 20 ppm) Target 2 (from Paul Hodgson) Stopped after small, but noticable deviations from nominal position noticed (next 2 slides) To be dismantled and examined for – Direction wrto stator of wear for each bearing – Systematic off-axis force due to coil quality or misalignment Achieved over 2,000,000 actuations - Record for target run in R78!
Feb. 10 T2 Acceleration and Position Left: acceleration vs. time Right: position vs. Time – ‘BCD’ is Beam Center Distance 22 Feb, 20103
4 Feb. 21 T2 Acceleration and Position Left: acceleration vs. time Right: position vs. Time – About 0.4 mm drop over 11 days 22 Feb, 20104
MICE User Run Feb. 18: took data sample for online reconstruction tests – 300 MeV/c π + with TOF1 trigger – Online Monitoring moved from DAQ machine Online Reconstruction computerin to reduce burden on DAQ machine – Successful test of online monitoring of live data from new location – Online monitoring reference plots, documentation being developed by Linda and Vassil Run (supposed to be easy) had problems which were solved – Remote neutron monitor wasn’t working Resolved by Linda, Pierrick, and James – TOF and KL voltages somehow got reset to low nominal setting (1141 V) Reset manually Need to keep and eye on this; remote possibility of cause by Pierrick working on EPICS software – ISIS changed timing from 40 Hz to 50 Hz which affected our target timing Timing worked out by Linda and Craig – Q4 power supply water flow meter tripped repeatedly, eventually turned off This one will be replaced this week and tested at full current New meters ordered 22 Feb, 20105
User Run Plans Reluctance to use up target pulse budget without working decay solenoid No runs planned, but requests in for Wednesday and Friday MICE data taking request with indefinite dates when decay solenoid working – Check and complete (emittance, momentum) matrix – Measure particle production vs. beam loss for different particle species – Proton absorber studies 22 Feb, 20106